# Solving and learning advective multiscale Darcian dynamics with the Neural Basis Method

### v1

@@ Methods
$\mathbf{q}+a\,\nabla p$ $=0\quad\text{in }\Omega,$ (M11)
$\nabla\cdot\mathbf{q}+\beta\,p$ $=0\quad\text{in }\Omega,$ (M12)
with mixed boundary conditions
$p=p_{D}\quad\text{on }\partial\Omega_{p},\qquad\mathbf{q}\cdot\mathbf{n}=q_{N}\quad\text{on }\partial\Omega_{q}.$
We define the residual components
$r_{m}$ $:=\tilde{\mathbf{q}}+a\nabla\tilde{p}$,
$r_{c}$ $:=\nabla\!\cdot\tilde{\mathbf{q}}+\beta\tilde{p}$,
$r_{D}$ $:=\tilde{p}-p_{D},$
$r_{N}$ $:=\tilde{\mathbf{q}}\cdot\mathbf{n}-q_{N},$
@@
The next lemma shows that this weighted residual norm is equivalent to the natural graph norm or error such that $\|R(\tilde{p},\tilde{\mathbf{q}})\|_{W}\asymp\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}}$. Consequently, we provide a computable residual–error certificate.
@@
$C_{1}\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}}\leq\|R(\tilde{p},\tilde{\mathbf{q}})\|_{W}\leq C_{2}\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}},$ (M13)
@@
The equivalence between $\|\cdot\|_{W}$ and the corresponding unweighted $L^{2}$-residual norm follows from $0<c_{\star}\leq w_{\star}\leq C_{\star}$. The error–residual equivalence (M13) is the $U$-ellipticity result for the first-order system least-squares operator associated with (M11)–(M12); see, e.g., [9, 22]. ∎
Residual–error certificate for the stabilized transport projection
We consider a single implicit transport substep $t^{m}\to t^{m+1}$ based on the upwind control-volume balance in (6). Let $c$ denote the exact concentration and let $\tilde{c}$ be its NBM approximation, and define the error $e_{c}:=\tilde{c}-c$ and the corresponding cell-wise time step-dependent residual $r_{\mathrm{up}}:=A_{\mathrm{up}}\tilde{c}-b$, so that $r_{\mathrm{up}}=A_{\mathrm{up}}e_{c}$. For an arbitrary cellwise scalar field $v=\{v_{K}\}$, define the pore-volume–scaled diagonal mass weight
$D:=\mathrm{diag}(\varepsilon|K|),\qquad\|v\|_{D}^{2}:=\sum_{K}(\varepsilon|K|)\,v_{K}^{2},$
and the upwind dissipation semi-norm over interior faces as follows. For each interior face $f=K|L$, let $\mathbf{n}_{K,f}$ be the unit normal on $f$ pointing outward from cell $K$, and set $\alpha_{f}:=\big|(\mathbf{u}\cdot\mathbf{n}_{K,f})_{f}\big|\,|f|.$ Then
$|v|_{\mathrm{up}}^{2}:=\frac{1}{2}\sum_{f=K|L}\alpha_{f}\,(v_{K}-v_{L})^{2}.$
We then define the transport energy norm
$\|v\|_{\mathcal{E},\mathrm{tr}}^{2}:=\frac{1}{\Delta t}\|v\|_{D}^{2}+|v|_{\mathrm{up}}^{2}.$
Since $|v|_{\mathrm{up}}^{2}\geq 0$, the transport energy norm of error controls the cellwise Euclidean error of concentration
$\|e_{c}\|_{2}\ \leq\ \sqrt{\frac{\Delta t}{\varepsilon|K|_{\min}}}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}},\qquad|K|_{\min}:=\min_{K}|K|.$ (T.3a)
@@
$C_{1}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}}\ \leq\ \|r_{\mathrm{up}}\|_{2}\ \leq\ C_{2}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}}.$ (M14)
@@
For the upwind control-volume operator $A=\frac{1}{\Delta t}D+V$, the symmetric part $V_{s}=\tfrac{1}{2}(V+V^{\top})$ induces the classical upwind jump dissipation, i.e., $v^{\top}V_{s}v=|v|_{\mathrm{up}}^{2}\geq 0$ for all cell vectors $v$; see, e.g., [7, 4]. Hence $G:=\frac{1}{\Delta t}D+V_{s}\succ 0$ and $\|v\|_{\mathcal{E},\mathrm{tr}}^{2}=v^{\top}Gv$. Writing $A=G+V_{k}$ with skew-symmetric $T_{k}=\tfrac{1}{2}(V-V^{\top})$, the DPG theory yields the residual–error equivalence in the induced dual norm, $\|e\|_{G}\asymp\|Ae\|_{G^{-1}}$; see, e.g., [19, 17, 18]. Finally, on a fixed discretization, the Euclidean residual $\|r\|_{2}$ is equivalent to $\|r\|_{G^{-1}}$ by standard spectral norm equivalence for symmetric positive definite matrices, which completes (M14). ∎
NBM–OL implementation
NBM–OL implementation is very structured as shown in Fig. 4. Below, we show a complete yet high-level description with the optional POD compression included (specific configurations are provide in SI.11). We begin by outlining the shared aspects of the learning pipeline used in the three experiments:
  • Case-I: permeability-varying Darcy flow operator;

### v2

@@ 4.1 Residual–error certificate for the weighted Darcy projection
$\mathbf{q}+a\,\nabla p$ $=0\quad\text{in }\Omega,$ (15)
$\nabla\cdot\mathbf{q}+\beta\,p$ $=0\quad\text{in }\Omega,$ (16)
with mixed boundary conditions $p=p_{D}$ on $\partial\Omega_{p}$ and $\mathbf{q}\cdot\mathbf{n}=q_{N}$ on $\partial\Omega_{q}$. We define the residuals
$\begin{array}[]{ll}r_{m}:=\tilde{\mathbf{q}}+a\nabla\tilde{p},&r_{c}:=\nabla\!\cdot\tilde{\mathbf{q}}+\beta\tilde{p},\\ r_{D}:=\tilde{p}-p_{D},&r_{N}:=\tilde{\mathbf{q}}\cdot\mathbf{n}-q_{N}.\end{array}$
@@
Then, Lemma 2 shows that this weighted residual norm is equivalent to the natural graph norm such that $\|R(\tilde{p},\tilde{\mathbf{q}})\|_{W}\asymp\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}}$. Consequently, we provide a computable residual–error certificate.
@@
$C_{1}\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}}\leq\|R(\tilde{p},\tilde{\mathbf{q}})\|_{W}\leq C_{2}\,\|(e_{p},e_{\mathbf{q}})\|_{\mathcal{E}},$ (17)
@@
The equivalence between $\|\cdot\|_{W}$ and the corresponding unweighted $L^{2}$-residual norm follows from $0<c_{\star}\leq w_{\star}\leq C_{\star}$. The error–residual equivalence (17) is the $U$-ellipticity result for the first-order system least-squares operator associated with (15)–(16); see, e.g., [9, 22]. ∎
4.2 Residual–error certificate for the stabilized transport projection
We next consider an implicit transport sub-timestep $t^{m}\to t^{m+1}$ based on the upwind control-volume balance in (11). Let $c$ denote the exact concentration and $\tilde{c}$ its NBM approximation, and define the error by $e_{c}:=\tilde{c}-c$. We further define the residual for this sub-timestep as $r_{\mathrm{up}}:=A_{\mathrm{up}}\tilde{c}-b$, so that $r_{\mathrm{up}}=A_{\mathrm{up}}e_{c}$. For an arbitrary cellwise scalar field $v=\{v_{K}\}$, we introduce the pore-volume scaled diagonal mass weight $D:=\mathrm{diag}(\varepsilon|K|),\,\|v\|_{D}^{2}:=\sum_{K}(\varepsilon|K|)\,v_{K}^{2}.$ For each interior face $f=K|L$, let $\mathbf{n}_{K,f}$ denote the unit normal on $f$ pointing outward from cell $K$, and set $\alpha_{f}:=\big|(\mathbf{u}\cdot\mathbf{n}_{K,f})_{f}\big|\,|f|.$ We can write the upwind dissipation semi-norm as $|v|_{\mathrm{up}}^{2}:=\frac{1}{2}\sum_{f=K|L}\alpha_{f}\,(v_{K}-v_{L})^{2}.$ We then define the transport energy norm $\|v\|_{\mathcal{E},\mathrm{tr}}^{2}:=\frac{1}{\Delta t}\|v\|_{D}^{2}+|v|_{\mathrm{up}}^{2}.$ Since $|v|_{\mathrm{up}}^{2}\geq 0$, the transport energy norm controls the cellwise Euclidean error of concentration
$\|e_{c}\|_{2}\ \leq\ \sqrt{\frac{\Delta t}{\varepsilon|K|_{\min}}}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}},\qquad|K|_{\min}:=\min_{K}|K|.$
@@ 4.2 Residual–error certificate for the stabilized transport projection
$C_{1}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}}\ \leq\ \|r_{\mathrm{up}}\|_{2}\ \leq\ C_{2}\,\|e_{c}\|_{\mathcal{E},\mathrm{tr}}.$ (18)
@@
For the upwind control-volume operator $A=\frac{1}{\Delta t}D+V$, the symmetric part $V_{s}=\tfrac{1}{2}(V+V^{\top})$ induces the classical upwind jump dissipation, i.e., $v^{\top}V_{s}v=|v|_{\mathrm{up}}^{2}\geq 0$ for all cell vectors $v$; see, e.g., [7, 4]. Hence $G:=\frac{1}{\Delta t}D+V_{s}\succ 0$ and $\|v\|_{\mathcal{E},\mathrm{tr}}^{2}=v^{\top}Gv$. Writing $A=G+V_{k}$ with skew-symmetric $T_{k}=\tfrac{1}{2}(V-V^{\top})$, the DPG theory yields the residual–error equivalence in the induced dual norm, $\|e\|_{G}\asymp\|Ae\|_{G^{-1}}$; see, e.g., [19, 17, 18]. Finally, on a fixed discretization, the Euclidean residual $\|r\|_{2}$ is equivalent to $\|r\|_{G^{-1}}$ by standard spectral norm equivalence for symmetric positive definite matrices, which completes (18). ∎